OpenStudy (anonymous):

Find the distance between P1(3, –195°) and P2(–4, –94°) on the polar plane. Round your answer to the nearest thousandth. I got 5.75 as my answer. Is this correct?

OpenStudy (anonymous):

\(\bbox [10pt, blue, border:5pt solid black]{\Huge\color{red}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!!!}}}\)

OpenStudy (anonymous):

Yes it is correct

OpenStudy (anonymous):

Revamped P1(3, –195°) x = r cos ∝ = 3 cos(-195) = 3(-0.966) x = -2.9 y = r sin α = 3 sin(-195) = 3(0.259) y = 0.8 P1(x, y) = (-2.9, 0.8) P2(–4, –94°) x = -4 cos(-94) = -4(-0.0698) = 0.28 y = -4 sin(-94) = -4(0.998) = -3.99 P2(0.28, -3.99) (-2.9, 0.8) and (0.28, -3.99) P1 to P2 distance is d = √(y2 – y1)² + (x2 – x1)² = √((-3.99 – 0.8)² + (0.28 + 2.9)²) = √(-4.79² + 3.18²) = √33.0565 = 5.75

OpenStudy (anonymous):

Thanks! Can you help me? @PyroYolka

OpenStudy (anonymous):

Awesome thank-you so much!

OpenStudy (anonymous):

P1(3, –195°) x = r cos ∝ = 3 cos(-195) = 3(-0.966) x = -2.9 y = r sin α = 3 sin(-195) = 3(0.259) y = 0.8 P1(x, y) = (-2.9, 0.8) P2(–4, –94°) x = -4 cos(-94) = -4(-0.0698) = 0.28 y = -4 sin(-94) = -4(0.998) = -3.99 P2(0.28, -3.99) (-2.9, 0.8) and (0.28, -3.99) P1 to P2 distance is d = √(y2 – y1)² + (x2 – x1)² = √((-3.99 – 0.8)² + (0.28 + 2.9)²) = √(-4.79² + 3.18²) = √33.0565 = 5.75

OpenStudy (anonymous):

hmmm....

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

@PyroYolka

OpenStudy (anonymous):

srry my os crashed close this question and open a new one :)